• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name B. Crowe

    Log entry time 00:01:39 on May 7,2006

    Entry number 173238

    keyword=Swing Shift Run Summary

    16:51 Started Production Run 4512. Replay (noticed a slight change in spectrum 18-N1 and Spectrum 17-V1 Left.
    18:08 Started Production Run 4513. Replay ok.
    18:33 Ended Run 2M counts
    18:35 Started Production Run 4514. Replay ok.
    19:34 Started Production Run 4515. Replay ok. Aq (56.611+-77.115)
    BPMA X= 0.471 Y = 1.255 BPMB X = 0.008 Y = -0.047
    20:35 Took NMR Measurement. Polarization 52.9%
    20:39 Started Production Run 4516. Replay ok. Aq (-47.783+-40.267)
    BPMA X =0.462 Y= 1.247 BPMB X = 0.015 Y = -0.064
    21:39 Started Production Run 4517. Replay (noticed double peak in total sum 22) BPMA X = 0.491 Y =1.238 BPMA X = 0.035 Y = -0.026
    Aq (76.054+-21.159)
    22:40 Started Production Run 4518. Replay ok. Big Bite HV tripped ended run early 869K counts. Beam experienced frequent trips due to the fast feed bac system. MCC worked on the probblem. In the meantime we took the target out of position and replaced it with the empty target position.
    23:42 MCC thinks that it was a magnet mismatch that caused the beam tripping problem. All is okay now . Started Production Run 4519.